# FADA: Fast Diffusion Avatar Synthesis with Mixed-Supervised Multi-CFG Distillation

### v1

@@ Abstract
Diffusion-based audio-driven talking avatar methods have recently gained attention for their high-fidelity, vivid, and expressive results. However, their slow inference speed limits practical applications. Despite the development of various distillation techniques for diffusion models, we found that naive diffusion distillation methods do not yield satisfactory results. Distilled models exhibit reduced robustness with open-set input images and a decreased correlation between audio and video compared to teacher models, undermining the advantages of diffusion models. To address this, we propose FADA (Fast Diffusion Avatar Synthesis with Mixed-Supervised Multi-CFG Distillation). We first designed a mixed-supervised loss to leverage data of varying quality and enhance the overall model capability as well as robustness. Additionally, we propose a multi-CFG distillation with learnable tokens to utilize the correlation between audio and reference image conditions, reducing the threefold inference runs caused by multi-CFG with acceptable quality degradation. Extensive experiments across multiple datasets show that FADA generates vivid videos comparable to recent diffusion model-based methods while achieving an NFE speedup of 4.17-12.5 times. Demos are available at our webpage http://fadavatar.github.io.
@@ 1 Introduction
Talking avatar synthesis aims to animate a given portrait image using driven video or audio. Due to the ease of obtaining audio inputs and their low usage barrier, there has been a surge of recent work in this area. Based on the technical route, talking avatar synthesis methods can be categorized into GAN-based[42, 29, 41, 39, 7, 17, 4], NeRF-based[5, 25, 34, 33, 36, 35], and the more recent diffusion model-based approaches[26, 1, 31, 2, 9, 12, 28]. Diffusion model-based methods, in particular, have gained attention due to their strong generative capabilities. These models exhibit excellent robustness in input images and produce vivid and highly expressive videos that naturally align with the audio. However, the current diffusion model-based talking head generation methods suffer from slow inference speeds, which hinder their practical application. Methods[9, 31, 1] based on Stable Diffusion (SD)[21], all require multiple denoising steps. To ensure the correlation between the audio input and the image input in the final video, multi-CFG (Classifier-Free Guidance) inference is often employed, further increasing the runtime. Although most methods achieve 30-40 denoising steps with the help of DDIM[23], they still face speed challenges due to CFG calculations
Recently, several works have explored diffusion distillation in text-to-image [22, 11, 19, 3, 13, 15, 24, 40] and text-to-video tasks [16] to accelerate inference. Nonetheless, there has been no direct exploration of diffusion distillation for talking avatar synthesis tasks. Our observations indicate that simply applying text-to-image and text-to-video diffusion distillation methods results in a noticeable decline in performance. Unlike text to image or video tasks, talking avatar synthesis involves two key conditions: audio and a reference image. Since audio has a relatively weaker influence on video generation[9, 26], the final output is jointly influenced by both conditions, necessitating careful parameter tuning in multi-CFG settings. The sensitivity to condition guidance often leads to severe artifacts when applying diffusion distillation methods with few-step inference.
@@
Based on the above considerations and designed methods, we implemented FADA, a diffusion distillation method for audio-driven talking head generation, achieving up to a 24$\times$ speed-up while generating vivid portrait videos comparable to recent undistilled diffusion models. In summary, the contributions of this paper include:
@@ 2.2 Diffusion Model Acceleration
Reducing the number of denoising runs is a direct way to accelerate diffusion models. SD-Turbo[22] leverages the ideas from GAN and introduces adversarial diffusion distillation to enhance the generation fidelity. SnapFusion[11] proposes progressive distillation to achieve fewer steps by learning the previous student model progressively. SwiftBrush[19] and SwiftBrushV2[3] utilizing variational score distillation to make it possible for texts-only distillation in text-to-image tasks. SDXL-lightning[13] mixes progressive distillation and adversarial distillation to achieve one-step image generation with a delicate training process. Applying consistency models [15, 24] to text-to-image tasks also helps reduce diffusion inference steps. LCM[15] views the guided reverse diffusion process as solving an augmented probability flow ODE. TCD[40] proposes trajectory consistency distillation to accurately trace the entire trajectory of the probability flow ODE in semi-linear form with an exponential integrator. However, there is a specially-designed distillation framework and techniques for talking avatar synthesis tasks.
@@ 3.1 Overall Framework and Preliminary
$L_{\mathrm{tea}}=\mathbb{E}_{z_{t},c,t,\epsilon}\left[\lVert\epsilon-\epsilon_{\mathrm{tea}}(z_{t},t,c;\theta_{\mathrm{tea}})\rVert_{2}^{2}\right]$ (1)
Where $c$ includes the audio feature, reference image and motion frames. Note that it is not necessary for the teacher model to be trained with flow-matching loss, and both $\epsilon$-prediction and $\upsilon$-prediction are valid for PeRFlow distillation. During the student distillation training, firstly same-interval $K$ time windows $\{[t_{k},t_{k-1})\}_{k=K}^{1}$ are created where $t_{k}=k/K,k\in\{0,1,...,K\}$. A time window $[t_{k},t_{k-1})$ is randomly sampled out, and the standpoint $z_{t_{k}}$ can be derived from the marginal distribution of the ground-truth targets, where $z_{t_{k}}=\sqrt{1-\sigma^{2}(t_{k})}z_{0}+\sigma(t_{k})\epsilon$ and $\sigma(t)$ denotes the noise schedule. Then the teacher-predicted ending point $\hat{z}_{t_{k-1}}$ of this time window can be predicted and solved by ODE solver $\Phi(\cdot)$:
$\hat{z}_{t_{k-1}}=\Phi(z_{t_{k}},t_{k},t_{k-1};\theta_{\mathrm{tea}})$ (2)
@@
$\hat{\epsilon}(t)=\frac{\hat{z}_{t_{k-1}}-\lambda_{k}z_{t_{k}}}{\eta_{k}}$ (4)
Where $\lambda_{k}=\sqrt{\alpha_{t_{k-1}}}/\sqrt{\alpha_{t_{k}}}$ and $\eta_{k}=\sqrt{1-\alpha_{t_{k-1}}}-\sqrt{1-\alpha_{t_{k}}}\sqrt{\alpha_{t_{k-1}}}/\sqrt{\alpha_{t_{k}}}$ respectively. Finally, the training objective of the distillation process can be obtained as:
$L_{\mathrm{distill}}=\mathbb{E}_{k,t\in[t_{k},t_{k-1}),z_{t},c,\epsilon}\left[\lVert\hat{\epsilon}(t)-\epsilon_{\mathrm{stu}}(\hat{z_{t}},t,c;\theta_{\mathrm{stu})}\rVert_{2}^{2}\right]$ (5)
@@ 3.2 Adaptive Mixed-Supervised Distillation
In our experiments, we found that distillation significantly degrades the model’s performance, especially overall video synthesis quality, as shown in Table 3. Increasing data can improve performance, but in audio-driven portrait generation, the weak control of audio over motion requires the model to learn high-quality motion patterns to fill in details not covered by audio. This necessitates stricter data filtering, reducing the available data.
Fortunately, this problem can be mitigated in our proposed distillation process. In our design, we first maintain the teacher training unchanged with only a high-quality dataset $\mathcal{A}$ to ensure its output stability. Then during the distillation stage, the data filtering strategies will be relaxed and some moderate-quality data samples ( a larger dataset $\mathcal{B}$ that includes $\mathcal{A}$ where $|\mathcal{B}|>|\mathcal{A}|$.) will be fed to student training. Since the student model mimics the teacher model’s output, it can maintain the high-quality generation capability of the teacher model even with the inclusion of lower-quality data. To further leverage the new data during the distillation stage, we introduce a ground-truth supervised loss $L_{\mathrm{gt}}$. The ground-truth noisy latent $z^{*}_{t}$ and ground-truth target noise $\epsilon^{*}(t)$ can be simply obtained by replacing $\hat{z}_{t_{k-1}}$ with ground truth ending point $z_{t_{k-1}}$ in Formula 3 and 4. Similar to distillation loss computation, the ground-truth loss $L_{\mathrm{GT}}$ and total loss $L_{\mathrm{total}}$ can be described as:
$L_{\mathrm{GT}}=\mathbb{E}_{k,t\in[t_{k},t_{k-1}),z_{t},c,\epsilon}\left[\lVert\epsilon^{*}(t)-\epsilon_{\mathrm{stu}}(z_{t}^{*},t,c;\theta_{\mathrm{stu})}\rVert_{2}^{2}\right]$ (6)
$L_{\mathrm{total}}=L_{\mathrm{distill}}+\mathcal{W}\cdot L_{\mathrm{GT}}$ (7)
@@
$\mathcal{W}=\mathcal{W}_{0}\cdot\left\{\begin{array}[]{lr}\mathcal{R}^{s},&\mathcal{R}<\mathcal{R}_{p}\\ \mathcal{R}_{p}^{s}(\mathcal{R}_{d}-\mathcal{R})/(\mathcal{R}_{d}-\mathcal{R}_{p}),&\mathcal{R}_{p}<=\mathcal{R}<\mathcal{R}_{d}\\ 0,&\mathcal{R}>=\mathcal{R}_{d}\end{array}\right.$ (8)
@@ 3.3 Multi-CFG Distillation with Learnable Token
In audio-driven talking avatar synthesis tasks, the Multi-CFG technique is utilized [9, 12] to achieve robust control of multiple conditions (reference image and audio) effectively,
Specifically, the single CFG calculates the direction from unconditional predicted values to a certain conditional predicted value and then moves a certain distance from the unconditional predicted values to obtain conditional predicted values. Furthermore, the Multi-CFG recursively computes the next CFG based on the first CFG. In this paper, multi-CFG inference result noise $\epsilon_{cfg}$ can be derived as follows[9]:
$\hat{\epsilon}_{cfg}=\mathrm{cfg}_{a}\times(\hat{\epsilon}_{a}-\hat{\epsilon}_{r})+\mathrm{cfg}_{r}\times(\hat{\epsilon}_{r}-\hat{\epsilon}_{b})+\hat{\epsilon}_{b}$ (9)
Where $\mathrm{cfg}_{a}$ and $\mathrm{cfg}_{r}$ indicate the CFG guidance scale of audio and reference condition. $\hat{\epsilon}_{a}$ refers to the noise prediction with both audio and reference conditions, while $\hat{\epsilon}_{r}$ removes the audio condition and $\hat{\epsilon}_{b}$ lacks audio and reference condition. Therefore, during inference, the model needs to run three times for each denoising step, which is time-consuming.
@@
$Emb_{\mathrm{cfg}}=\mathrm{cfg}_{a}\times(\gamma_{a}-\gamma_{r})+\mathrm{cfg}_{r}\times(\gamma_{r}-\gamma_{b})+\gamma_{b}$ (10)
@@ 4.1 Experimental Settings
The number of time windows in PeRFlow[32] is 4. All results of FADA are conducted with 6-step inferences, where the two time windows near $T=1$ contain two inference steps while the others contain one inference step.
@@ 4.2 Comparison with State-of-the-Arts
Sadtalker[37] - 2.953 8.765 171.8 36.64 2.248 3.435 7.870 24.93 25.35 1.559
Hallo[31] 80 3.505 9.079 53.99 35.96 2.426 3.922 7.917 21.71 20.15 1.337
EchoMimic[1] 60 3.307 10.37 54.71 35.37 3.018 3.994 9.391 18.71 19.01 1.328
V-Express[28] 50 2.946 9.415 117.8 65.09 2.414 3.482 7.382 48.03 30.91 1.506
Ours-Balanced 18 3.588 8.998 50.36 33.67 2.202 3.951 7.925 13.32 18.39 1.362
Ours-Fast 6 3.550 8.746 54.69 35.01 2.604 3.927 7.996 16.67 18.51 1.635
@@
For the purpose of quantitative comparisons, we evaluate the generated talking avatar videos with several metrics. We assess the image quality with IQA[30] metrics and evaluate the audio-visual synchronization distance with a pre-trained SyncNet[20] as Sync-D. FID and FVD[27] (with 3D ResNet) are assessed to evaluate the image/video- level fidelity difference between generated results and ground truths in the image inception space. E-FID (Expression-FID)[26] measures the inception distance between the generated and ground truths in expression parameter space. NFE-D indicates the number of function evaluations of the denoising net. For instance, a 25-step DDIM inference with 3 CFG runs will lead to 75 NFE-D.
Since our proposed FADA is a diffusion distillation framework, we mainly compare recent open-source diffusion-based state-of-the-art methods in talking portrait generation, including V-Hallo[31], EchoMimic[1], and V-Express[28]. Additionally, for reference, we also include the state-of-the-art GAN-based method SadTalker[37].
@@
As shown in Table 2, we conduct quantitative comparison experiments with CelebV-HQ and HDTF test sets. Ours-Balanced refers to our proposed method with mixed-supervised distillation with PeRFlow loss. Based on it, Ours-Fast utilizes multi-CFG distillation to achieve about 3-time speedup and gets a slight performance dropping. Ours-Balanced achieves better IQA, FVD, FID, and E-FID on the challenging open-set CelebVHQ test set compared to the baselines. It also leads or is comparable in most metrics on the simpler HDTF test set. Ours-Fast maintains similar performance in IQA and Sync-D metrics with a slight drop in FVD, FID, and E-FID, which is still comparable to the baselines in most metrics. In terms of audio synchronization, the trend is similar. Notably, VExpress results show almost only lip movements, which, despite good metrics, do not produce satisfactory video quality. Finally, in terms of speedup, it is evident that both Ours-Balanced and Ours-Fast achieve significant speedup using the same SD base model as the baselines.
@@
As illustrated in Figure 2 and in the supplementary video materials, Ours-Fast and Ours-Balanced maintain good video quality and vivid audio-visual expressiveness. We highly recommend readers watch the supplementary videos since it is difficult for the static image to show the audio-motion synchronization and temporal consistency. V-Express shows color differences and inferior identity-preserving abilities in all cases. Sadtalker generates good identity-preserving videos at the image level, but the head motions and lip movements are not well-synced with the emotion and pace of the given audio. EchoMimic and Hallo generate acceptable expressions but there are temporal jitters that damage the visual quality. Moreover, by observing case-4 and case-6, we can find that FADA produces more full and plump pronunciations than baselines.

### v2

@@ Abstract
Diffusion-based audio-driven talking avatar methods have recently gained attention for their high-fidelity, vivid, and expressive results. However, their slow inference speed limits practical applications. Despite the development of various distillation techniques for diffusion models, we found that naive diffusion distillation methods do not yield satisfactory results. Distilled models exhibit reduced robustness with open-set input images and a decreased correlation between audio and video compared to teacher models, undermining the advantages of diffusion models. To address this, we propose FADA (Fast Diffusion Avatar Synthesis with Mixed-Supervised Multi-CFG Distillation). We first designed a mixed-supervised loss to leverage data of varying quality and enhance the overall model capability as well as robustness. Additionally, we propose a multi-CFG distillation with learnable tokens to utilize the correlation between audio and reference image conditions, reducing the threefold inference runs caused by multi-CFG with acceptable quality degradation. Extensive experiments across multiple datasets show that FADA generates vivid videos comparable to recent diffusion model-based methods while achieving an NFE speedup of 4.17-12.5 times. Demos are available at our webpage https://fadavatar.github.io.
@@ 1 Introduction
Talking avatar synthesis aims to animate a given portrait image using driven video or audio. Due to the ease of obtaining audio inputs and their low usage barrier, there has been a surge of recent work in this area.
Based on the technical route, talking avatar synthesis methods can be categorized into GAN-based[48, 34, 47, 44, 8, 19, 4], NeRF-based[5, 30, 39, 38, 41, 40], and the more recent diffusion model-based approaches[31, 1, 36, 2, 10, 14, 33]. Diffusion model-based methods, in particular, have gained attention due to their strong generative capabilities. These models exhibit excellent robustness in input images and produce vivid and highly expressive videos that naturally align with the audio.
However, the current diffusion model-based talking head generation methods suffer from slow inference speeds, which hinder their practical application. Methods[10, 36, 1] based on Stable Diffusion (SD)[23], all require multiple denoising steps. Moreover, to ensure the correlation between the audio input and the image input in the final video, multi-CFG (Classifier-Free Guidance) inference is often employed, further increasing the runtime. Although most methods achieve 30-40 denoising steps with the help of DDIM[27], they still face speed challenges due to CFG calculations.
Recently, several works have explored diffusion distillation in text-to-image [24, 12, 21, 3, 15, 17, 28, 46] and text-to-video tasks [18] to accelerate inference. Nonetheless, there has been no direct exploration of diffusion distillation for talking avatar synthesis tasks. Our observations indicate that simply applying text-to-image and text-to-video diffusion distillation methods results in a noticeable decline in performance.
Unlike text to image or video tasks, talking avatar synthesis involves two key conditions: audio and a reference image. Since audio has a relatively weaker influence on video generation[10, 31], the final output is jointly influenced by both conditions, necessitating careful parameter tuning in multi-CFG settings. The sensitivity to condition guidance often leads to severe artifacts when applying diffusion distillation methods with few-step inference.
@@
Based on the above considerations and designed methods, we implemented FADA, a diffusion distillation method for audio-driven talking head generation, achieving up to a 12.5$\times$ speed-up while generating vivid portrait videos comparable to recent undistilled diffusion models. In summary, the contributions of this paper include:
@@ 2.2 Diffusion Model Acceleration
Reducing the number of denoising runs is a direct way to accelerate diffusion models. SD-Turbo[24] leverages the ideas from GAN and introduces adversarial diffusion distillation to enhance the generation fidelity. SnapFusion[12] proposes progressive distillation to achieve fewer steps by learning the previous student model progressively. SwiftBrush[21] and SwiftBrushV2[3] utilizing variational score distillation to make it possible for texts-only distillation in text-to-image tasks. SDXL-lightning[15] mixes progressive distillation and adversarial distillation to achieve one-step image generation with a delicate training process. Applying consistency models [17, 28] to text-to-image tasks also helps reduce diffusion inference steps. LCM[17] views the guided reverse diffusion process as solving an augmented probability flow ODE. TCD[46] proposes trajectory consistency distillation to accurately trace the entire trajectory of the probability flow ODE in semi-linear form with an exponential integrator. However, there is no specially-designed distillation framework and techniques for talking avatar synthesis tasks to our knowledge.
@@ 3.1 Overall Framework and Preliminary
$L_{\mathrm{tea}}=\mathbb{E}_{z_{t},c,t,\epsilon}\left[\lVert\epsilon-\epsilon_{\mathrm{tea}}(z_{t},t,c;\theta_{\mathrm{tea}})\rVert_{2}^{2}\right],$ (1)
where $c$ includes the audio feature, reference image and motion frames.
Note that it is not necessary for the teacher model to be trained with flow-matching loss, and both $\epsilon$-prediction and $\upsilon$-prediction are valid for PeRFlow distillation. During the student distillation training, firstly same-interval $K$ time windows $\{[t_{k},t_{k-1})\}_{k=K}^{1}$ are created where $t_{k}=k/K,k\in\{0,1,...,K\}$. A time window $[t_{k},t_{k-1})$ is randomly sampled out, and the standpoint $z_{t_{k}}$ can be derived from the marginal distribution of the ground-truth targets, where $z_{t_{k}}=\sqrt{1-\sigma^{2}(t_{k})}z_{0}+\sigma(t_{k})\epsilon$ and $\sigma(t)$ denotes the noise schedule. Then the teacher-predicted ending point $\hat{z}_{t_{k-1}}$ of this time window can be predicted and solved by ODE solver $\Phi(\cdot)$:
$\hat{z}_{t_{k-1}}=\Phi(z_{t_{k}},t_{k},t_{k-1};\theta_{\mathrm{tea}}),$ (2)
@@
$\hat{\epsilon}(t)=\frac{\hat{z}_{t_{k-1}}-\lambda_{k}z_{t_{k}}}{\eta_{k}},$ (4)
Where $\lambda_{k}=\sqrt{\alpha_{t_{k-1}}}/\sqrt{\alpha_{t_{k}}}$ and $\eta_{k}=\sqrt{1-\alpha_{t_{k-1}}}-\sqrt{1-\alpha_{t_{k}}}\sqrt{\alpha_{t_{k-1}}}/\sqrt{\alpha_{t_{k}}}$ respectively. Finally, the training objective of the distillation process can be obtained as
$L_{\mathrm{distill}}=\mathbb{E}_{k,t\in[t_{k},t_{k-1}),z_{t},c,\epsilon}\left[\lVert\hat{\epsilon}(t)-\epsilon_{\mathrm{stu}}(\hat{z_{t}},t,c;\theta_{\mathrm{stu})}\rVert_{2}^{2}\right],$ (5)
@@ 3.2 Adaptive Mixed-Supervised Distillation
In our experiments, we found that distillation significantly degrades the model’s performance, especially overall video synthesis quality, as shown in Table 3. Increasing data can improve performance, but in audio-driven portrait generation, the weak control of audio over motion requires the model to learn high-quality motion patterns to fill in details not covered by audio. This necessitates stricter data filtering, reducing the available data quantity.
Fortunately, this problem can be mitigated in our proposed distillation process. In our design, we first maintain the teacher training unchanged with only a high-quality dataset $\mathcal{A}$ to ensure its output stability. Then during the distillation stage, the data filtering strategies will be relaxed and some moderate-quality data samples ( a larger dataset $\mathcal{B}$ that includes $\mathcal{A}$ where $|\mathcal{B}|>|\mathcal{A}|$.) will be fed to student training. Since the student model mimics the teacher model’s output, it can maintain the high-quality generation capability of the teacher model even with the inclusion of lower-quality data. To further leverage the new data during the distillation stage, we introduce a ground-truth supervised loss $L_{\mathrm{gt}}$. The ground-truth noisy latent $z^{*}_{t}$ and ground-truth target noise $\epsilon^{*}(t)$ can be simply obtained by replacing $\hat{z}_{t_{k-1}}$ with ground truth ending point $z_{t_{k-1}}$ in Formula 3 and 4. Similar to distillation loss computation, the ground-truth loss $L_{\mathrm{GT}}$ and total mixed loss $L_{\mathrm{total}}$ can be described as
$L_{\mathrm{GT}}=\mathbb{E}_{k,t\in[t_{k},t_{k-1}),z_{t},c,\epsilon}\left[\lVert\epsilon^{*}(t)-\epsilon_{\mathrm{stu}}(z_{t}^{*},t,c;\theta_{\mathrm{stu})}\rVert_{2}^{2}\right],$ (6)
$L_{\mathrm{total}}=L_{\mathrm{distill}}+\mathcal{W}\cdot L_{\mathrm{GT}},$ (7)
@@
$\mathcal{W}=\mathcal{W}_{0}\cdot\left\{\begin{array}[]{lr}\mathcal{R}^{s},&\mathcal{R}\in[0,\mathcal{R}_{p})\\ \mathcal{R}_{p}^{s}(\mathcal{R}_{d}-\mathcal{R})/(\mathcal{R}_{d}-\mathcal{R}_{p}),&\mathcal{R}\in[\mathcal{R}_{p},\mathcal{R}_{d})\\ 0,&\mathcal{R}\in[\mathcal{R}_{d},+\infty)\end{array}\right.$ (8)
@@ 3.3 Multi-CFG Distillation with Learnable Token
In audio-driven talking avatar synthesis tasks, the Multi-CFG technique is utilized [10, 14] to achieve robust control of multiple conditions (reference image and audio) effectively, Specifically, the single CFG calculates the direction from unconditional predicted values to a certain conditional predicted value and then moves a certain distance from the unconditional predicted values to obtain conditional predicted values. Furthermore, the Multi-CFG recursively computes the next CFG based on the first CFG. In this paper, multi-CFG inference result noise $\epsilon_{cfg}$ can be derived as follows[10]:
$\hat{\epsilon}_{cfg}=\mathrm{cfg}_{a}\times(\hat{\epsilon}_{a}-\hat{\epsilon}_{r})+\mathrm{cfg}_{r}\times(\hat{\epsilon}_{r}-\hat{\epsilon}_{b})+\hat{\epsilon}_{b},$ (9)
where $\mathrm{cfg}_{a}$ and $\mathrm{cfg}_{r}$ indicate the CFG guidance scale of audio and reference condition. $\hat{\epsilon}_{a}$ refers to the noise prediction with both audio and reference conditions, while $\hat{\epsilon}_{r}$ removes the audio condition and $\hat{\epsilon}_{b}$ lacks both audio and reference.
Therefore, during inference, the model needs to run three times for each denoising step, which is time-consuming.
@@
$Emb_{\mathrm{cfg}}=\mathrm{cfg}_{a}\times(\gamma_{a}-\gamma_{r})+\mathrm{cfg}_{r}\times(\gamma_{r}-\gamma_{b})+\gamma_{b}.$ (10)
@@ 4.1 Experimental Settings
The number of time windows in PeRFlow[37] is 4. All results of FADA are conducted with 6-step inferences, where the two time-windows near $T=1$ contain two inference steps while the others contain one inference step.
@@ 4.2 Comparison with State-of-the-Arts
Sadtalker[42] - $2.953^{\pm.086}$ 8.765 171.8 36.64 2.248 $3.435^{\pm.069}$ 7.870 24.93 25.35 1.559
Hallo[36] 80 $3.505^{\pm.091}$ 9.079 53.99 35.96 2.426 $3.922^{\pm.079}$ 7.917 21.71 20.15 1.337
EchoMimic[1] 60 $3.307^{\pm.104}$ 10.37 54.71 35.37 3.018 $3.994^{\pm.083}$ 9.391 18.71 19.01 1.328
V-Express[33] 50 $2.946^{\pm.123}$ 9.415 117.8 65.09 2.414 $3.482^{\pm.071}$ 7.382 48.03 30.91 1.506
Ours-Balanced 18 $3.588^{\pm.091}$ 8.998 50.36 33.67 2.202 $3.951^{\pm.082}$ 7.925 13.32 18.39 1.362
Ours-Fast 6 $3.550^{\pm.093}$ 8.746 54.69 35.01 2.604 $3.927^{\pm.081}$ 7.996 16.67 18.51 1.635
@@
For the purpose of quantitative comparisons, we evaluate the generated talking avatar videos with several metrics. We assess the image quality with IQA[35] metrics and evaluate the audio-visual synchronization distance with a pre-trained SyncNet[22] as Sync-D. FID and FVD[32] (with 3D ResNet) are assessed to evaluate the image/video-level fidelity difference between generated results and ground truths in the image inception space. E-FID (Expression-FID)[31] measures the inception distance between the generated and ground truths in expression parameter space. NFE-D indicates the number of function evaluations of the denoising net. For instance, our teacher model conducts 25-step DDIM inference with 3 CFG runs, leading to 75 NFE-D.
Since our proposed FADA is a diffusion distillation framework, we mainly compare recent open-source diffusion-based state-of-the-art methods in talking portrait generation, including Hallo[36], EchoMimic[1], and V-Express[33]. Additionally, for reference, we also include the state-of-the-art GAN-based method SadTalker[42].
@@
[…]
@@
As illustrated in Figure 2 and in the supplementary video materials, Ours-Fast and Ours-Balanced maintain good video quality and vivid audio-visual expressiveness. We highly recommend readers watch the supplementary videos since it is difficult for the static image to show the audio-motion synchronization and temporal consistency. V-Express shows color differences and inferior identity-preserving abilities in all cases. Sadtalker generates good identity-preserving videos at the image level, but the head motions and lip movements are not well-synced with the emotion and pace of the given audio. EchoMimic and Hallo generate acceptable expressions, but there are temporal jitters that damage the visual quality. Moreover, by observing case-4 and case-6, we can find that FADA produces more full and plump pronunciations than baselines.